# Accelerating the Exploration of Thermal Materials via a Synergistic Strategy of DFT and Interpretable Deep Learning

### v9

@@ II.2 Deepening physical interpretability through sensitivity analysis
Figure 3: Feature correlation and sensitivity analysis. (a) Heatmap of the correlation matrix. The numbers in each cell represent the Pearson correlation coefficients between the corresponding features on the horizontal and vertical axes. The color gradient visually reflects the pairwise correlation strength. (b)-(e) Different sensitivity analysis methods are applied to evaluate feature importance for both the KAN and SISSO models.
When constructing the dataset, we curated a selection of closely related features, such as heat capacity at constant volume ($C_{\rm v}$) and heat capacity at constant pressure ($C_{\rm p}$), as well as isothermal bulk modulus ($K_{\rm i}$) and static (adiabatic) bulk modulus ($K_{\rm s}$). The purpose of this approach was to ensure that the ML models achieve reliable accuracy. The correlation coefficient heatmap can reveal the potential degree of linear association between features, as illustrated in Figure 3(a). It is important to clarify that $K_{\rm i}$ and $K_{\rm s}$ are not entirely identical concepts [52, 53, 54]. Both are temperature-dependent, and the difference between them increases as the temperature rises, but due to the minimal difference at 300K, a near-linear relationship appears, which is also observed between the two specific heat capacities [52]. Interestingly, in addition to the previously mentioned bulk moduli and heat capacities, the Debye temperature $\Theta$, vibrational free energy per atom $F$, and vibrational entropy per atom ${S_{\rm{v}}}$ also show a high degree of linear correlation. The quantitative relationship between vibrational free energy and vibrational entropy is defined by the following equation [55]:
$F=Q-T{S_{\rm{v}}}$ (2)
@@
$+\frac{{\hbar\left\langle{{\omega^{2}}}\right\rangle\left({\vec{\sigma}}\right)}}{{24{k_{\rm{B}}}T}}-\frac{{{\hbar^{4}}\left\langle{{\omega^{4}}}\right\rangle\left({\vec{\sigma}}\right)}}{{2880k_{\rm{B}}^{3}{T^{3}}}}+\cdots$ (3)
[…]
$F\approx\left\langle{\ln\left(\omega\right)}\right\rangle\left(\sigma\right){k_{\rm{B}}}T$ (4)
@@
As shown in Figure 3, we calculated the mean absolute values of various sensitivity indices, which provide the global values for SHAP and LIME. It is important to note that, regardless of the white-box machine learning surrogate model used, results obtained under the assumption of feature independence lack persuasiveness. In Figure 3(b), significant differences exist among the results from different methods, making it difficult to determine which is more reliable. In Figure 3(d), although different analysis methods yield similar feature sensitivities, correlated features should exhibit similar sensitivities [64]. Evidently, the conclusions from Figure 3(b) and (d) do not satisfy this premise. Therefore, conclusions drawn from sensitivity analysis methods that ignore feature interaction effects lack credibility.
The sensitivity analysis results in Figure 3(c) and (e) are relatively reasonable and satisfy the premise of ”similar sensitivities for correlated features,” but differences still exist. For instance, in Figure 3(e), the total effect indices of $C_{\rm p}$ and $C_{\rm v}$ are higher than those of other features, yet their first-order effect indices are approximately zero. A total effect index slightly higher than the first-order effect index indicates that the independent contributions of the features themselves are minimal, but their combinations with other features exert additional influence on the model output. In the symbolic SISSO model (Eq. (1)), specific heat capacities appear in the form $\left|{{C_{\rm v}}-{C_{\rm p}}}\right|$. Such interaction terms reduce the contribution of individual specific heat variables but result in a higher total contribution. However, a first-order effect index of zero suggests that $\log(\kappa_{\rm L})$ is entirely independent of specific heat capacities.
@@
When considering empirical models like Slack’s, the sensitivity analysis results of the SISSO symbolic model are contradictory and unsustainable, as they underestimate the influence of specific heat capacities on $\log(\kappa_{\rm L})$. Furthermore, despite the presence of other highly correlated variables, the removal of redundant features causes a noticeable performance degradation in the SISSO model, as shown in Figure 4(a). This indicates that SISSO fails to accurately capture the correlations between physical quantities and exhibits poor robustness, whereas the KAN model shows the opposite behavior. For a primary feature, a ${K_{\rm{i}}}$ value exceeding ${{K}^{\text{T}}}$ indicates a strong correlation with other features, while ${{K}^{\text{T}}}=0$ signifies an exact correlation with other inputs [28, 65]. From this perspective, KAN is more effective than SISSO at overcoming feature dependency. This advantage can be attributed to KAN’s neural network architecture. As illustrated in Figure 4(b), in the absence of pruning, the neurons in KAN’s hidden layers are fully connected to both preceding and subsequent layers. This design allows each feature to independently and equitably contribute to $\log(\kappa_{\rm L})$. Simultaneously, KAN leverages backpropagation to update the weights of various nodes, ensuring the model achieves optimal accuracy. This process not only guarantees fair participation of all features in model training but also effectively distinguishes their relative contribution.
Ultimately, we retained $C_{\rm v}$, $\alpha$, $F$, and $K_{\rm s}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while enabling fast and accurate predictions via CGCNN, thereby facilitating high-throughput screening of new materials. As shown in Figure 4(a), the four selected features are sufficient to describe the physical mechanism of $\kappa_{\rm L}$ with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. In contrast to KAN, SISSO is characterized by a relatively limited number of operators (see Eq. (12)), and its descriptors typically have a dimensionality of no more than 3. Exceeding this limit would lead to much greater computational resource demands compared to KAN and black-box models. These complexity constraints limit SISSO’s expressive capacity and hinder its ability to capture complex physical feature-mapping relationships.
$+1.4{{\rm{e}}^{-19.72d_{3}^{2}}}-0.48{{\rm{e}}^{-63.21d_{4}^{2}}}$ (5)
where
$\quad+1.56{e^{-13.65{{\left({-\alpha-0.17}\right)}^{2}}}}+0.85{{\rm{e}}^{-26.37{{\left({-{C_{\rm v}}-0.11}\right)}^{2}}}}$
$\quad-0.86\cosh\left({5.35{C_{\rm v}}-1.75}\right)+2.31$
$\quad+0.38{\mathop{\rm atan}\nolimits}\left({8.56{K_{\rm s}}-0.63}\right)$
$d_{3}$ $={\left({0.38-{K_{\rm s}}}\right)^{3}}-0.09\sin\left({6.2{C_{\rm v}}-6.16}\right)$
$\quad-0.08{\mathop{\rm atan}\nolimits}\left({4.8F-1.66}\right)+0.47$
$\quad-0.26{{\rm{e}}^{-38.44{{\left({-\alpha-0.06}\right)}^{2}}}}$
$\quad+0.28{e^{-2.53{{\left({-F-0.02}\right)}^{2}}}}$
$\log{\left(\kappa_{\rm L}\right)_{{\rm{SISSO}}}}=$ $1.53+0.0079\frac{{F\log\left({{C_{\rm{v}}}}\right)}}{{\sqrt[3]{{{C_{\rm{v}}}}}}}$
$-45.57\sqrt[3]{{{\alpha^{2}}{C_{\rm{v}}}{K_{\rm{s}}}}}$ (6)
Following feature reduction, the $R^{2}$ of the analytical expression derived by KAN was 0.9639, exhibiting minimal degradation in accuracy compared to its pre-reduction value of 0.9655. In contrast, as illustrated in Figure 4(a), SISSO experienced a pronounced decline in accuracy. Consequently, we assert that KAN more effectively captures the mapping relationship between $\kappa_{\rm L}$ and the features, while also demonstrating exceptional robustness.
Figure 5: Correlation between the 4 key primary features and $\kappa_{\rm L}$. (a)-(d) Two-dimensional mappings of different features with $\log(\kappa_{\rm L})$, where the feature values have been min-max normalized. The closed solid lines represent contour lines of the sample density distribution, while the dashed lines indicate fifth-order polynomial fitting curves. (e) The mapping between $\log(\kappa_{\rm L})$ and $F$, $K_{\text{s}}$, with the color bar indicating the $\log(\kappa_{\rm L})$ of each sample. The dashed line represents the mean $\log(\kappa_{\rm L})$ value of 1 for the samples.
Figure 6: Predicting key features and qualitatively assessing $\kappa_{\rm L}$ via CGCNN. (a), (b), and (c) display scatter plots comparing the CGCNN-predicted values of three features with the DFT-calculated values on their respective test sets, where $R^{2}_{F}=0.940$, $R^{2}_{K_{\rm s}}=0.980$ and $R^{2}_{\alpha}=0.910$.
II.3 Screening potential thermal insulators/conductors
Unlike simple descriptors that can be directly generated by programs, such as those based on composition (e.g., Magpie [66]) or structure (e.g., Coulomb matrix [67]), the physical features involved in the aforementioned reduced dataset are relatively complex and not easily constructed physical quantities. Inspired by previous studies [68, 69], our approach is to use a two-stage prediction method. We predict primary physical features by CGCNNs, and then use the predicted primary features to perform high-throughput predictions of the $\kappa_{\rm L}$ for unlabeled materials. CGCNNs have achieved accuracy comparable to or even surpassing DFT in predicting energy-related and mechanical properties [34]. In our application, CGCNNs delivered fairly accurate results for the prediction of all primary features except for $C_{\text{v}}$, where the prediction accuracy was notably lower (see Supplementary Information for details).Therefore, we focus on first screening a subset of potential thermal insulating/conductive materials using primary features that can be reliably predicted qualitatively, and then calculating $C_{\text{v}}$ using DFT. This approach is more efficient than directly using DFT or MD methods to calculate $\kappa_{\rm L}$.
Figure 7: $k$-means clustering and t-SNE 2D visualization. (a) Labeled training samples from AFLOW are clustered into four groups, each represented by a different color. (b) The same scatter plot as in (a), but the samples are colored according to their $\log(\kappa_{\rm L})$, with black pentagrams representing unlabeled candidates from the MP database. (c) and (d) are zoomed-in views showing the thermal insulator/thermal conductor candidates in the t-SNE space.
@@ II.3 Screening potential thermal insulators/conductors
mp-1184997 KRbAu2 -105.1144 14.1935 $2.36\times 10^{-4}$ 36.4862 0.2169
mp-1097263 Cs2RbNa -144.966 16.8264 $4.90\times 10^{-4}$ / /
mp-10378 Cs3Sb -143.5 11.2367 $2.42\times 10^{-4}$ / /
mp-1097633 Cs2KRb -145.791 14.89677 $5.10\times 10^{-4}$ / /
mp-635413 Cs3Bi -168.671 13.74844 $2.50\times 10^{-4}$ / /
Based on the information from Figure 3(a) and Figure 5(a)-(d), it can be concluded that $\kappa_{\rm L}$ is strongly and positively linearly correlated with both $F$ and $K_{\rm s}$. Materials with higher F and $K_{\text{s}}$ values generally exhibit higher $\kappa_{\rm L}$, a relationship that is further confirmed in Figure 5(e). Therefore, the preliminary qualitative screening of thermal insulators and conductors based on $F$ and $K_{\text{s}}$ is highly beneficial for accelerating the exploration of new materials. The predictions of $F$, $K_{\text{s}}$ and $\alpha$ using CGCNN are shown in Figure 6(a)-(c), with $R^{2}$ values reaching 0.94, 0.98 and 0.91, respectively. We extracted a total of 2,246 samples from the Materials Project (MP) [70] with the following selection criteria:
@@
  • Number of atoms (nsites) $<$ 5;
@@
For this purpose, we first attempt to use $k$-means clustering [71] on samples from the AFLOW database (e.g., the labeled dataset used for model training in Section II.1) to evaluate its effectiveness. Since reliable quantitative estimation is currently only feasible for the three primary features $F$, $K_{\rm s}$, and $\alpha$, we retain only these three features for each sample during clustering. In $k$-means clustering, the optimal number of clusters is typically determined using the elbow method [72] or the silhouette score [73]. However, in our experiments, these two methods yielded inconsistent results, making it unclear which one is more convincing (see Supplementary Information). Fortunately, through manual inspection, we found that when $K=4$, the clustering results already exhibit a clear and meaningful trend.
@@
In Figure 7(a), all training samples are distinctly clustered into seven clusters, $C_{1}$ to $C_{4}$, with well-defined boundaries between clusters. Mapping the values of $\log(\kappa_{\rm L})$ onto the classification results in Figure 7(a), as shown in Figure 7(b), we observe that as t-SNE component 1 increases, $\log(\kappa_{\rm L})$ exhibits an upward trend. This allows us to identify $C_{1}$ as the cluster corresponding to thermal insulators and $C_{4}$ as the cluster representing thermal conductors. More importantly, the clustering results indicate that $\log(\kappa_{\rm L})$ can be accurately and qualitatively assessed using only the three features $F$, $K_{\rm s}$, and $\alpha$. Based on the labeled dataset, we obtained a CGCNN pre-trained model for the initial features and applied it to predict $F$, $K_{\rm s}$, and $\alpha$ for unlabeled samples from the MP database. These predicted features were then clustered in the same manner as the labeled samples. In Figure 7(b), we selected the 10 most promising thermal insulator/conductor candidates from clusters $C_{1}$ and $C_{4}$ (see Table 3). The zoomed-in views are shown in Figures 7(c) and (d), where these candidates are found in the same clusters as the labeled samples with the lowest/highest $\log(\kappa_{\rm L})$, respectively. Experimental results proved that only 4 out of the 10 candidates (CN2, CrB4, Be2C and KRbAu2) effectively utilize DFT and MD to calculate $C_{\rm v}$ and $\kappa_{\rm L}$, respectively. The subsequent validation process will also focus on these 4 feasible structures.
Figure 8: Validation of the KAN predictions. For (a) Be2C, (b) CN2 and (c) CrB4, a relaxation time of 5 ns is sufficient for the $\kappa_{\rm L}$ to converge. The deep red dashed line represents the $\kappa_{\rm L}$ predicted by the KAN model, while the light red band indicates the prediction range with a ±5% error margin. The $\kappa_{\rm L}$ calculated by MD is represented by the blue line, illustrating its convergence behavior with respect to relaxation time. (d) The blue circular scatter points and connecting lines represent the temperature dependence curve of $\kappa_{\rm L}$ for KRbAu2, while the red dashed line represents the $\kappa_{\rm L}$ predicted by the KAN model. (e) shows a histogram of the $\kappa_{\rm L}$ distribution in the training dataset.

### v10

@@ III.2 Deepening physical interpretability through sensitivity analysis
Figure 4: Correlation matrix scatterplot displaying pairwise relationships among key thermal and mechanical properties of materials. Diagonal panels show histograms of individual variable distributions. Off-diagonal scatterplots are color-coded by $\kappa_{\rm L}$ on a logarithmic scale, highlighting how $\kappa_{\rm L}$ varies with respect to each feature pair. Pearson correlation coefficients ($R$) quantify linear relationships between variables, aiding interpretation of interdependencies relevant for thermal transport analysis.
Figure 5: Feature sensitivities of the KAN and SISSO surrogate models. (a) and (b) are based on the KAN model, while (c) and (d) correspond to the SISSO model. Specifically, (a) and (c) show results obtained using SHAP, LIME, and Sobol indices, whereas (b) and (d) display sensitivities calculated using the Kucherenko index. The $y$-axis indicates the sensitivity value, reflecting the relative importance of each feature.
When constructing the dataset, we curated a selection of closely related features, such as heat capacity at constant volume ($C_{\rm v}$) and heat capacity at constant pressure ($C_{\rm p}$), as well as isothermal bulk modulus ($K_{\rm i}$) and static (adiabatic) bulk modulus ($K_{\rm s}$). The purpose of this approach was to ensure that the ML models achieve reliable accuracy. The correlation coefficient heatmap can reveal the potential degree of linear association between features, as illustrated in Fig. 4. It is important to clarify that $K_{\rm i}$ and $K_{\rm s}$ are not entirely identical concepts [83, 84, 85]. Both are temperature-dependent, and the difference between them increases as the temperature rises, but due to the minimal difference at 300K, a near-linear relationship appears, which is also observed between the two specific heat capacities [83]. Interestingly, in addition to the previously mentioned bulk moduli and heat capacities, the Debye temperature $\Theta$, vibrational free energy per atom $F$, and vibrational entropy per atom $S_{\rm{v}}$} also show a high degree of linear correlation. The quantitative relationship between vibrational free energy and vibrational entropy is defined by the following equation [86]:
$F=Q-T{S_{\rm{v}}},$ (18)
@@
$+\frac{{\hbar\left\langle{{\omega^{2}}}\right\rangle\left({\vec{\sigma}}\right)}}{{24{k_{\rm{B}}}T}}-\frac{{{\hbar^{4}}\left\langle{{\omega^{4}}}\right\rangle\left({\vec{\sigma}}\right)}}{{2880k_{\rm{B}}^{3}{T^{3}}}}+\cdots,$ (19)
where the $\vec{\sigma}$ denotes the configuration of $A$ and $B$ atoms on the lattice, while $\left\langle\cdot\right\rangle$ represents the average operation (per atom) over the Brillouin zone. $E_{0}$} refers to the fully relaxed ground state energy, $\omega$ is the vibrational frequency of a phonon mode, $k_{\rm{B}}$} and $\hbar$ correspond to the Boltzmann and Planck constants, respectively. As the temperature approaches the Debye temperature, setting $T=\Theta$, we have:
$F\approx\left\langle{\ln\left(\omega\right)}\right\rangle\left(\sigma\right){k_{\rm{B}}}T,$ (20)
@@
The assumption in Fig. 5(a) and (c) is that SHAP, LIME, and Sobol indices treat the input features as independent, which leads to qualitatively incorrect conclusions. This issue arises because Fig. 4 shows that the feature groups $\{F,S_{\rm v},\Theta\}$, $\{C_{\rm p},C_{\rm v}\}$, and $\{K_{\rm s},K_{\rm i}\}$ exhibit high mutual correlations. Considering that strongly correlated features are expected to exhibit similar sensitivities in the decision-making process [93], the independence assumption becomes problematic. A more specific issue can be observed in Fig. 5(c), where, due to the dimensional constraint of the SISSO model (see Eq. (17)), only features $\{\Theta,\gamma,C_{\rm p},C_{\rm v},K_{\rm i},S_{\rm v}\}$ are retained and involved in the model construction. Even under this premise, both $S_{i}$ and $S_{i}^{\rm T}$ for $\{C_{\rm p},C_{\rm v}\}$ are exactly zero with respect to the remaining excluded variables. This is attributed to the presence of the term $\left|{{C_{\rm{p}}}-{C_{\rm{v}}}}\right|$ in Eq. (17). From a global perspective, $C_{\rm p}$ and $C_{\rm v}$ exhibit nearly synchronous variations, resulting in an almost invariant value for their difference, which highlights a fundamental limitation of global sensitivity indices. In contrast, the local sensitivity metrics, such as SHAP and LIME, yield more reasonable interpretations. However, a remaining issue is that the sensitivity of $\gamma$ is evaluated to be zero—reflected in the LIME index where it shows sensitivity comparable to those of excluded variables. While the cause of this phenomenon remains unclear, it is evident that sensitivity analysis that neglects feature dependence yields qualitatively incorrect results for both KAN and SISSO-based surrogate models.
@@
By contrast, the conclusions drawn from Kucherenko indices appear far more physically plausible, though the choice of surrogate model still introduces subtle differences. In Fig. 5(c), the individual contributions of features are quantified by the first-order index $K_{1}$. When KAN is used as the surrogate, all strongly correlated physical variables are assigned nearly identical sensitivity values, owing to the superior ability of Kucherenko indices to capture inter-feature dependencies. Moreover, the vanishing $K_{1}$ for $\rm SG$} aligns well with physical intuition, since existing models of $\kappa_{\rm L}$ do not explicitly incorporate space group symmetry [94, 95, 96]. Interestingly, all features exhibit similar total sensitivity $K^{\rm T}$, which stems from the fully connected architecture of KAN: each feature contributes equally to the modelling of $\log(\kappa_{\rm L})$, rendering them structurally and interactively equivalent [40, 28].
Fig. 5(d) shows the outcome when SISSO is used as the surrogate. Compared to Fig. 5(c), the improvement lies in the ability of Kucherenko indices to assign similar sensitivity values to correlated but unmodelled inputs. However, due to the inherent limitations of the SISSO model, the contributions of $\{\gamma,C_{\rm v},C_{\rm p}\}$ remain underestimated. Taken together, only the combination of KAN and Kucherenko indices yields sensitivity results consistent with physical prior knowledge in our case.
@@
mp-1184997 KRbAu2 105.1144-105.1144 14.1935 $2.36\times 10^{-4}$ 36.4862 0.2169
mp-1097263 Cs2RbNa 144.966-144.966 16.8264 $4.90\times 10^{-4}$ / /
mp-10378 Cs3Sb 143.5-143.5 11.2367 $2.42\times 10^{-4}$ / /
mp-1097633 Cs2KRb 145.791-145.791 14.89677 $5.10\times 10^{-4}$ / /
mp-635413 Cs3Bi 168.671-168.671 13.74844 $2.50\times 10^{-4}$ / /
Ultimately, we retained $C_{\rm v}$, $\alpha$, $F$, and $K_{\rm s}$ as the features for the reduced dataset, as these features preserve as much of the original information as possible while enabling fast and accurate predictions via CGCNN, thereby facilitating high-throughput screening of new materials. As shown in Fig. 6(a), the four selected features are sufficient to describe the physical mechanism of $\kappa_{\rm L}$ with a high degree of confidence. Similar to other black-box models, KAN exhibits only a negligible decline in performance with the reduction in feature count. In contrast, SISSO’s performance deteriorates significantly. This decline is likely due to the complexity constraints of the SISSO model. In contrast to KAN, SISSO is characterized by a relatively limited number of operators (see Eq. (6)), and its descriptors typically have a dimensionality of no more than 3. Exceeding this limit would lead to much greater computational resource demands compared to KAN and black-box models. These complexity constraints limit SISSO’s expressive capacity and hinder its ability to capture complex physical feature-mapping relationships.
Figure 7: Partial dependency plots between each feature and $\kappa_{\rm L}$. In the diagonal subplots, the horizontal axis corresponds to the values of the feature shown at the bottom, while the red vertical axis represents the values of $\kappa_{\rm L}$. In the off-diagonal subplots, each scatter point’s color and shape correspond to different clusters defined in $k$-means clustering, and the background color indicates the estimated thermal conductivity $\mathbb{E}_{\mathcal{X}_{C}}\left[\hat{f}(\mathcal{X}_{S}|\mathcal{X}_{C})\right]$. Here, $\mathbb{E}_{\mathcal{X}_{C}}\left[\hat{f}(\mathcal{X}_{S}|\mathcal{X}_{C})\right]$ refers to the predicted $\kappa_{\rm L}$ value when only the selected features $\mathcal{X}_{S}$ are varied, while all other features $\mathcal{X}_{C}$ are fixed at their dataset-wide mean values. For example, if $\mathcal{X}_{S}=\{F,K_{\rm s}\}$, then $\mathcal{X}_{C}=\{C_{\rm v},\alpha\}$. The estimation $\mathbb{E}_{\mathcal{X}_{C}}\left[\hat{f}(\mathcal{X}_{S}|\mathcal{X}_{C})\right]$ is intended to infer the $\kappa_{\rm L}$ in a virtual space where real samples are absent, based on the KAN-derived symbolic model.
$+1.4{{\rm{e}}^{-19.72d_{3}^{2}}}-0.48{{\rm{e}}^{-63.21d_{4}^{2}}},$ (21)
where
$\quad+1.56{e^{-13.65{{\left({-\alpha-0.17}\right)}^{2}}}}+0.85{{\rm{e}}^{-26.37{{\left({-{C_{\rm v}}-0.11}\right)}^{2}}}},$
$\quad-0.86\cosh\left({5.35{C_{\rm v}}-1.75}\right)+2.31$
$\quad+0.38{\mathop{\rm atan}\nolimits}\left({8.56{K_{\rm s}}-0.63}\right),$
$d_{3}$ $={\left({0.38-{K_{\rm s}}}\right)^{3}}-0.09\sin\left({6.2{C_{\rm v}}-6.16}\right)+0.47$
$\quad-0.08{\mathop{\rm atan}\nolimits}\left({4.8F-1.66}\right)-0.26{{\rm{e}}^{-38.44{{\left({-\alpha-0.06}\right)}^{2}}}},$
$\quad+0.28{e^{-2.53{{\left({-F-0.02}\right)}^{2}}}}.$
$\log{\left(\kappa_{\rm L}\right)_{{\rm{SISSO}}}}=1.53+0.0079\frac{{F\log\left({{C_{\rm{v}}}}\right)}}{{\sqrt[3]{{{C_{\rm{v}}}}}}}-45.57\sqrt[3]{{{\alpha^{2}}{C_{\rm{v}}}{K_{\rm{s}}}}}.$ (22)
Figure 8: $k$-means clustering and $t$-SNE 2D visualization. (a) Labeled training samples from AFLOW are colored according to their $\kappa_{\rm L}$, with white pentagrams representing unlabeled candidates from the MP database. (b) The same scatter plot as in (a), but the samples are clustered into 4 groups, each represented by a different color and shape. (c) and (d) are zoomed-in views showing the thermal insulator/thermal conductor candidates in the $t$-SNE space. The $x$-axis and $y$-axis represent the two components of the $t$-SNE embedding, respectively.
Following feature reduction, the $R^{2}$ of the analytical expression derived by KAN was 0.9639, exhibiting minimal degradation in accuracy compared to its pre-reduction value of 0.9655. In contrast, as illustrated in Fig. 6(a), SISSO experienced a pronounced decline in accuracy. Consequently, we assert that KAN more effectively captures the mapping relationship between $\kappa_{\rm L}$ and the features, while also demonstrating exceptional robustness.
Figure 9: Predicting key features and qualitatively assessing $\kappa_{\rm L}$ via CGCNN. (a), (b), and (c) display parity plots comparing the CGCNN-predicted values of three features with the DFT-calculated values on their respective test sets, where $R^{2}_{F}=0.940$, $R^{2}_{K_{\rm s}}=0.980$, $R^{2}_{\alpha}=0.910$, and $R^{2}_{C_{\rm v}}=0.528$.
III.3 Screening potential thermal insulators/conductors
Unlike straightforward descriptors derived from composition (e.g., Magpie [97]) or structure (e.g., Coulomb matrix [98]), the physical features in the reduced dataset are more intricate and not readily computable from fundamental physical quantities. Inspired by previous studies [99, 100], our approach is to use a two-stage prediction method. We predict primary physical features by CGCNNs, and then use the predicted primary features to perform high-throughput predictions of the $\kappa_{\rm L}$ for unlabeled materials. CGCNNs have achieved accuracy comparable to or even surpassing DFT in predicting energy-related and mechanical properties [35]. In our application, CGCNNs delivered fairly accurate results for the prediction of all primary features except for $C_{\rm v}$, where the prediction accuracy was notably lower (see Supplementary Information for details).Therefore, we focus on first screening a subset of potential thermal insulating/conductive materials using primary features that can be reliably predicted qualitatively, and then calculating $C_{\rm v}$ using DFT. This approach is more efficient than directly using DFT or MD methods to calculate $\kappa_{\rm L}$.
[…]
@@ III.3 Screening potential thermal insulators/conductors
  • Number of atoms (nsites) << 5;
@@
For this purpose, we first attempt to use $k$-means clustering [102] on samples from the AFLOW database (e.g., the labeled dataset used for model training in Section III.1) to evaluate its effectiveness. Since reliable quantitative estimation is currently only feasible for the three primary features $F$, $K_{\rm s}$, and $\alpha$, we retain only these three features for each sample during clustering. In $k$-means clustering, the optimal number of clusters is typically determined using the elbow method [103] or the silhouette score [104]. However, in our experiments, these two methods yielded inconsistent results, making it unclear which one is more convincing (see Supplementary Information). Fortunately, through manual inspection, we found that when $K=4$, the clustering results already exhibit a clear and meaningful trend.
Figure 10: MD and DFT validation of candidates’ $\kappa_{\rm L}$. For (a) Be2C, (b) CN2, and (c) CrB4, the MD results are shown as blue points with error bars, together with the running average curve (orange). The deep red dashed line denotes the $\hat{\kappa}_{\rm L}$ predicted by the KAN model, and the shaded red band indicates the $\pm 5\%$ interval. The MD results demonstrate that a relaxation time of 5 ns is sufficient for convergence of $\kappa_{\rm L}$. (d) For KRbAu2, the blue circles connected by lines show the temperature dependence of $\kappa_{\rm L}$ obtained from DFT, while the deep red dashed line and shaded band indicate the KAN prediction and its $\pm 5\%$ error interval.
@@
In Fig. 8(a), all training samples are distinctly clustered into seven clusters, $C_{1}$ to $C_{4}$, with well-defined boundaries between clusters. Mapping the values of $\log(\kappa_{\rm L})$ onto the classification results in Fig. 8(a), as shown in Fig. 8(b), we observe that as $t$-SNE component 1 increases, $\log(\kappa_{\rm L})$ exhibits an upward trend. This allows us to identify $C_{1}$ as the cluster corresponding to thermal insulators and $C_{4}$ as the cluster representing thermal conductors. More importantly, the clustering results indicate that $\log(\kappa_{\rm L})$ can be accurately and qualitatively assessed using only the three features $F$, $K_{\rm s}$, and $\alpha$. Based on the labeled dataset, we obtained a CGCNN pre-trained model for the initial features and applied it to predict $F$, $K_{\rm s}$, and $\alpha$ for unlabeled samples from the MP database. These predicted features were then clustered in the same manner as the labeled samples. In Fig. 8(b), we selected the 10 most promising thermal insulator/conductor candidates from clusters $C_{1}$ and $C_{4}$ (see Table 3). The zoomed-in views are shown in Figures 8(c) and (d), where these candidates are found in the same clusters as the labeled samples with the lowest/highest $\log(\kappa_{\rm L})$, respectively. Experimental results proved that only 4 out of the 10 candidates (CN2, CrB4, Be2C and KRbAu2) effectively utilize DFT and MD to calculate $C_{\rm v}$ and $\kappa_{\rm L}$, respectively. The subsequent validation process will also focus on these 4 feasible structures.